### v1

# Dualtronics: leveraging both faces of polar semiconductors

Abstract
Unlike non-polar semiconductors such as silicon, the broken inversion symmetry of the wide bandgap semiconductor gallium nitride leads to a large electronic polarization along a unique crystal axis. This makes the two surfaces of the semiconductor wafer perpendicular to the polar axis dramatically different in their physical and chemical properties. In the last three decades, the cation (gallium) face of gallium nitride has been used for photonic devices such as LEDs and lasers. Though the cation face has also been predominantly used for electronic devices, the anion (nitrogen) face has recently shown promise for high electron mobility transistors due to favorable polarization discontinuities. In this work we show that it is possible to make photonic devices on the cation face, and electronic devices on the anion face, of the same semiconductor wafer. This opens the possibility for dualtronics in polar semiconductors, where electronic, photonic, and acoustic properties can be implemented on opposite faces of the same wafer, dramatically enhancing the functional capabilities of this revolutionary semiconductor family.
Single crystals of semiconductors are macroscopic molecules with periodic repeating basis units[1]. In a perfect semiconductor crystal wafer, the atomic location and type at one point in space repeats itself hundreds of microns from one surface to the other, and hundreds of millimeters from one edge to the other. In non-polar semiconductors such as silicon, the crystal orientation is chosen carefully to maximize electronic performance of devices on the surface. Electronic devices such as transistors are made only on one (say the top) surface of silicon and not the other. Even if the other surface was used, the parallel surfaces of a cubic crystal are identical by symmetry. This means flipping a silicon wafer upside down does not change the chemical or electronic properties of the top surface. Thus, heterostructures and electronic devices made on the two surfaces in an identical manner exhibit identical properties.
Figure 1: Schematic of plasma-assisted molecular beam epitaxial growth of the HEMT-LED. The gray arrows indicate the chronological order of the growth procedure. (a) Ga-bonding of the GaN substrate and MBE growth of the (In,Ga,Al)N LED along the metal polar direction. (b) Unloading, cleaning, flipping, Ga-bonding and reloading of the sample. (c) MBE growth of the GaN/AlGaN HEMT along the nitrogen polar direction.
The wide bandgap semiconductors Gallium Nitride (GaN) and Aluminum Nitride (AlN) have a wurtzite crystal structure with an underlying hexagonal lattice [2]. This crystal structure breaks inversion symmetry along the [0001] orientation (the c-axis). The two surfaces perpendicular to the c-axis of a single crystal wafer of these polar semiconductors therefore exhibit dramatically different physical properties: flipping the crystal wafer is analogous to flipping a bar magnet. The chemical properties of the two sides is so distinct that they are used to identify metal or nitrogen surface polarity. While the metal-polar surface is inert to most chemicals, the N-polar surface of both GaN and AlN etches vigorously in solutions with bases, such as KOH and TMAH, or acids such as H3PO4 [3, 4].
Dramatic differences in electronic properties of heterostructures are observed for the two faces [5, 6]. For example, if a $\sim$10 nm thin coherently strained epitaxial GaN layer is deposited on the N-polar surface of AlN, a two-dimensional (2D) electron gas is formed at the GaN/AlN heterojunction quantum well due to the discontinuity of the conduction band combined with the discontinuity of the electronic polarization across the heterojunction [7]. But when an identical $\sim$10 nm coherently strained epitaxial GaN layer is deposited on the metal-polar surface of AlN, a 2D hole gas is formed at the heterojunction quantum well [8]. These polarization-induced, electrically conductive 2D electron and hole gases are formed in crystals nominally free of chemical impurities such as donor or acceptor dopants. (Al,Ga)N/GaN high electron mobility transistors (HEMTs) using such polarization-induced conductive channels demonstrate outstanding performance in high-power and high-speed applications [9, 10, 11, 12].
To date, only a single face of c-axis oriented GaN single crystal wafers is used for either photonic, or electronic devices [13, 14, 15]. In this work, the two polarities on the opposite faces of a GaN single crystal are combined to realize a photonic device on one side, and an electronic device on the other. A heterostructure quantum well on the N-polar side is used to generate a polarization-induced high mobility 2D electron gas (2DEG), and a quantum well (In,Ga,Al)N pn diode heterostructure is realized on the opposite metal-polar side. This two-sided wafer is then processed, first into HEMTs on the N-polar side, followed by blue quantum well light emitting diodes (LEDs) on the metal-polar side. Successful operation of the HEMT and LED devices is observed, allowing the switching and modulation of blue LEDs by HEMTs on the other side of the wafer.
Figure 1 shows how dual side epitaxy for dualtronics was achieved. The starting wafer is a high transparency bulk n-type O-doped GaN substrate grown by ammonothermal method [16, 17] with mobile electron concentration $\sim$1018 cm-3, and threading dislocation density $\sim$104 cm-2 with both Ga-polar and N-polar sides chemo-mechanically polished to obtain atomically smooth surfaces. The N-polar surface of the GaN substrate was gallium-bonded to a GaN/sapphire carrier wafer as shown in Figure 1(a). Because the surface bonded to the carrier wafer needs to be used again, we chose gallium bonding to the GaN surface of a GaN/sapphire carrier wafer to ensure desired elements (Ga, N) at this bonding interface. Using molecular beam epitaxy (MBE), we first grew a blue (In,Ga,Al)N LED structure consisting of a quantum well active region inside a pn heterojunction on the Ga-polar face of GaN. The wafer was then detached from the carrier wafer and flipped as shown in Figure 1(b). The Ga was removed from the N-polar side, and the metal-polar p-type GaN surface of the LED side was gallium-bonded on to a clean GaN/sapphire carrier wafer. Subsequently, a N-polar GaN/AlGaN/GaN heterostructure was grown on the N-polar face again by MBE to create a high-electron mobility 2DEG to serve as the conduction channel of the HEMT as shown in Figure 1(c). The chosen growth conditions for the LED and HEMT structures are described in the methods section.
Figure 2: (a) Schematic of the HEMT-LED heterostructures grown on both faces of a single crystal c-plane n-GaN substrate. (b) HAADF-STEM image showing the GaN/Al0.40Ga0.60N/GaN HEMT and (c) atomic resolution image corresponding to the uppermost GaN/Al0.40Ga0.60N heterojunction interface which hosts the 2DEG. (d) iDPC image in the uppermost GaN layer of the HEMT indicating that the nitrogen polarity follows that of the substrate to the surface. (e) HAADF-STEM image showing the LED quantum well, electron blocking layer, cladding and contact layers and (f) atomic resolution image corresponding to the LED In0.07Ga0.93N/In0.17Ga0.83N/In0.07Ga0.93N single quantum well. (g) iDPC image of the p-InGaN contact layers of the LED indicating the metal polarity follows that of the substrate.
Figure 2(a) shows the precise layer thicknesses and compositions of the HEMT on the N-polar side on the top and the LED on the metal-polar side on the bottom of the single GaN wafer achieved after the double side epitaxy. Figures 2(b-g) show the corresponding scanning transmission electron microscopy (STEM) images.
The MBE grown HEMT on the N-polar side started with the growth of a 500 nm thick unintentionally doped (UID) GaN buffer layer as shown in Figure 2(a), followed by 5 nm GaN:Si and 5 nm Al0.40Ga0.60N:Si layers with a Si concentration of 3 $\times$ $10^{18}\mathrm{cm}^{-3}$ capped with undoped 15 nm Al0.40Ga0.60N and 10 nm GaN. Silicon doping below the 15 nm AlGaN layer was employed to prevent the formation of a polarization-induced 2D hole gas at the bottom N-polar AlGaN/GaN interface[18]. The bulk energy bandgap of GaN is $E_{g1}=3.4$ eV, and of Al0.40Ga0.60N is $E_{g2}=4.5$ eV. For the N-polar HEMT the high-angle annular dark-field (HAADF) STEM image in Figure 2(b) shows the 20 nm AlGaN layer near the surface with sharp interfaces with GaN on both sides. Figure 2(c) zooms in to show the lattice image of the top GaN/AlGaN heterojunction where the 2DEG resides on mostly the GaN side. A sharp heterointerface between GaN and AlGaN is observed with the atomic planes clearly resolved. Figure 2(d) shows an integrated differential phase contrast (iDPC) image of the GaN HEMT channel layer, proving the N-polar crystal structure along the growth direction: the N-atoms located vertically below the brighter Ga atoms.
[…]
The MBE grown LED heterostructure on the metal-polar side seen on the bottom side of Figure 2(a) consists of a 150 nm GaN:Si buffer layer followed by a 50 nm In0.07Ga0.93N/40 nm In0.07Ga0.93N active region with a single 2.5 nm thick In0.17Ga0.83N quantum well embedded inbetween.
This photonic active region was followed by p-type doped layers starting with a 20 nm Al0.09Ga0.91N electron blocking layer with Mg concentration of 3 $\times$ $10^{19}\mathrm{cm}^{-3}$ followed by a 150 nm thick GaN layer with a Mg concentration of 6 $\times$ $10^{18}\mathrm{cm}^{-3}$. This LED heterostructure was capped with p-InGaN contact layers consisting of 40 nm thick In0.02Ga0.98N and 5 nm thick In0.15Ga0.85N doped with Mg at the level of 1 $\times$ $10^{20}\mathrm{cm}^{-3}$ and 5 $\times$ $10^{20}\mathrm{cm}^{-3}$, respectively, to form low resistance ohmic contacts for hole injection into the LED active region, while simultaneously to be transparent to the light emitted from the active region. Figure 2(e) shows a HAADF-STEM image of each of these LED layers indicating a uniform quantum well and the surrounding active region along the plane of the wafer. The atomically resolved image in Figure 2(f) shows the 2.5 nm thick In0.17Ga0.83N quantum well and the InGaN cladding layers. The group III element fractions and doping profiles for both the LED and HEMT structure are presented in the time-of-flight secondary ion mass spectrometry (ToF-SIMS) data shown in Extended Data Fig. 1, which indicate that the measured values are as intended. Importantly, there is no noticeable interdiffusion of dopants or group-III broadening of the metal-polar LED layers, indicating interface and doping control is not affected by the second growth step.
Figure 3: (a)-(d) Device processing flow for the double-sided HEMT-LED. Starting from the as-grown heterostructures, the gray arrows follow the independent processing steps chronologically, with the metal-polar LED being processed after the N-polar HEMT. (e) A three-dimensional representation of the complete device. (f) Optical microscope images of the as-processed sample, focused on the LEDs (left) and the HEMTs (right). The N-polar HEMTs are oriented upwards, forming the uppermost surface. For scale, the diameter of the large LED anode contact is 140 µm. (g) Scanning electron microscope (SEM) images of the HEMTs on the N-polar GaN surface (left) and the LEDs on the metal-polar GaN surface (right).
Figure 2(g) shows an iDPC image of the p-InGaN contact layers on the metal-polar LED side. The LED is indeed metal-polar viewed in the growth direction with the substrate out-of-view above. This confirms that the polarity of the crystal substrate locks the polarity of the epitaxial layers, and that the crystalline registry goes through the entirety of the GaN wafer over hundreds of microns. This is why the 2DEG forms on the the N-polar side of the wafer, while the metal-polarity is retained throughout the active regions, quantum well, and electron blocking and contact layers of the LED. The high crystalline perfection of the LED heterostructure also proves that the growth of the HEMT heterostructure on the N-polar side at a high temperature and with nitrogen plasma does not destroy or degrade the LED structure during the rather harsh thermal and chemical conditions. The surface of the as-grown heterostructures after the completion of both epitaxial steps as shown in Extended Data Fig. 2 exhibit sub-nanometer roughness atomic steps characteristic of step-flow growth, indicating the successful dual side epitaxy. X-ray diffraction measurements indicate high crystallinity, sharp heterointerfaces, and coherently strained layers to the GaN substrate seen in Extended Data Fig. 3.
Figure 4: Device characteristics of the HEMTs and LEDs, operating independently. (a) Drain current (black, solid) and transconductance (gray, solid) of an N-polar HEMT as a function of gate-source voltage, operating at a drain-source voltage of 5 V. The black, dashed line indicates a channel sheet current-density of 1 A/mm. (b) Linear plot showing the family of curves for a HEMT (solid black) for a gate-source voltage ranging from 1.75 (on) to -3.25 V (off). On the same scale, the linear current-voltage characteristics of a 400 µm diameter LED is shown as well (blue, solid). The dimensions of the measured HEMTs are LSD = 4 µm, LG = 1.5 µm, and WG = 50 µm. (c) Semi-log plots showing the drain current (black, solid) and gate current (black, dashed) versus gate-source voltage, for a drain-source voltage of 5 V. Here the transistor current corresponds to the left vertical axis. The horizontal, black, dashed line indicates a channel sheet current-density of 1 A/mm. Similarly, corresponding to the right vertical axis is the LED current (blue, solid) as a function of forward bias for a 400 µm diameter device. (d) Electroluminescence spectra of a metal-polar, 400 µm diameter LED. The injection current density ranges from 1 to 100 A/cm2. The top-right inset shows a camera image of the sample with an LED in the on-state.
For successful dualtronics, the survival of the thin LED heterostructure layers during the epitaxial growth conditions of the HEMT side needs to be replicated for the steps of device fabrication. The N-polar surface is particularly difficult to protect chemically because it reacts strongly with several solutions that are typically used for surface cleaning and conditioning during the processing steps. For example, the natural etching of the N-polar surface of GaN in solutions of KOH, TMAH, or H3PO4 form pyramidal facets that significantly roughen the pristine, flat surface [22]. This property of surface roughening is put to good use for efficient light extraction in blue LEDs. But such roughening would destroy the N-polar HEMT desired here, which requires sub-nanometer smoothness to preserve the high mobility 2DEG and for the fabrication of source, drain, and gate electrodes. Figure 3 shows the procedure adopted for the fabrication of the HEMT and the LEDs on the opposite sides of the same wafer, along with the final dualtronic device. A more detailed overview schematic of the fabrication of the dualtronic device is shown in Extended Data Fig. 4.
Figure 5: (a) Energy-band diagrams of the HEMT and LED indicating the off and on-states. (b) Circuit schematic of the monolithic HEMT-LED, taking into account the back-gating effect of the conductive GaN substrate. (c)-(d) Monolithic switching measurements, modulating between on-off (c) and between bright and dim modes (d). The gate-source modulation voltage is shown in red, the photodiode voltage while modulating the LED in solid blue, and the background (LED off state) photodiode voltage in dashed blue.
Figure 3(a) shows the starting layer structure which is first processed into a HEMT on the N-polar side shown in Figure 3(b). Ti/Au low resistance source and drain ohmic contacts are formed on regrown GaN layers that are heavily n-type doped with Silicon. A thin gate SiO2/Al2O3 dielectric layer was deposited followed by a Ni/Au gate metal stack to complete the HEMT fabrication. The HEMT layer was then protected by a photoresist coating and flipped upside down to the metal-polar side up for LED fabrication for each photolithography step.
Figure 3(c) shows that for the fabricated LED, Pd/Au metallization was used for the p-type anode contact. The high work function metal Palladium aligns the metal Fermi level close to the valence band holes of the p-type doped InGaN cap layer of the LED side. The thin p-InGaN reduces the contact resistance by enhancing the free hole concentration and by lifting the valence band energy-maximum. To form the n-type cathode contact, mesa regions were etched by removing the entire active region and terminating the etch on the buried n-type GaN layer to reveal a flat surface. Ti/Au metal stacks were deposited on this flat n-GaN surface to form a low resistance contact between electrons at the Fermi surface of the metal and the electrons in the GaN conduction band.
Figure 3(d) shows the completed dualtronic structure after removal of the protective photoresist coating, and Figure 3(e) shows a cross-section of it flipped upside down. This schematic emphasizes the single-crystal nature of the whole structure, across the bulk of the wafer indicated by the wavy break in the thick GaN layer between the HEMT on the top and the LED on the bottom. Figure 3(f) shows an optical microscope image focused on the top HEMT surface on the left, and the LED focused on the right. The energy bandgap of the bulk wafer between the HEMT and LED mesas is $E_{g1}=3.4$ eV of GaN, which is optically transparent to visible light, allowing for imaging of the LEDs through the wafer from the HEMT side as seen in Figure 3(f). This also emphasizes that the transistor layers are transparent to visible wavelengths, and are therefore attractive for dualtronic integration with LEDs as transparent thin-film transistors (TFTs) but, as seen later, with dramatically higher performance than existing TFTs made of oxide or organic semiconductors. Figure 3(g) shows the scanning electron microscope (SEM) images of the HEMTs on the N-polar GaN surface (left) and the LEDs on the metal-polar GaN surface (right).
Figure 4(a) shows the measured transfer characteristic and transconductance of the N-polar HEMT. A threshold voltage of $V_{T}\approx-3$ V is measured with an on-current drive exceeding 1 A/mm at a drain voltage of $V_{DS}=5$ V and gate voltage $V_{GS}\approx 3$ V. The transistor transconductance $g_{m}=\partial I_{D}/\partial V_{GS}$ peaks at $V_{GS}\approx-1.6$ V. Figure 4(b) shows the measured $I_{D}-V_{DS}$ output characteristics of the HEMT for various gate voltages overlaid with the LED current-voltage characteristics. The transistor delivers enough output current to drive the LED with variable drive currents controlled by the gate voltage up to 50 mA or more. Figure 4(c) shows the LED current and the HEMT current for $V_{DS}=5$ V in the logarithmic scale indicating that the HEMT can control the LED current and hence its optical output intensity over several orders of magnitude. Figure 4(d) shows the measured output spectra of the LED for injection current densities ranging from 1 to 100 A/cm2. The emission intensity increases with the injection current density, and the peak emission wavelength blue-shifts from $\lambda\approx 470$ nm at 1 A/cm2 to $\lambda\approx 450$ nm at 100 A/cm2 in the blue regime of the visible spectrum. The blueshift is caused by screening of the internal polarization fields which result in a quantum confined Stark effect (QCSE), as well as the Burstein-Moss effect. The inset in Figure 4(d) shows the bright blue electroluminescent glow of the blue LED slightly below the center of the wafer on the metal-polar side. The output and transfer characteristics of the HEMTs before and after the fabrication of the LEDs is shown in the Extended Data Fig. 5. It was confirmed that after the LED processing, there was minimal degradation of the HEMT with a less then 0.3 V threshold voltage shift and negligible change in the output and gate leakage current. This minimal degradation highlights the feasibility of using double-side epitaxy and processing for reliable heterogeneous device integration on both crystal faces.
Figure 5(a) shows the layer structures and corresponding energy band diagrams of the HEMT on the left and the LED on the right when they are connected in the circuit format shown in Figure 5(b). The source of the HEMT is connected to the anode of the LED to inject current to light up the LED. The light emitted from this dualtronic device is measured with a photodiode. A customized probe setup was fabricated to probe both HEMTs and LEDs without the need to flip the sample, shown schematically in Extended Data Fig 6. When the gate voltage of the HEMT is below the transistor threshold ($V_{GS}<V_{T}$), the triangular quantum well is lifted above the Fermi level in the energy band diagram of the HEMT as shown, and the transistor is off. The corresponding energy band diagram of the LED is shown on the right. When the HEMT gate voltage is above threshold, the bottom of the triangular quantum well is pulled below the Fermi level as seen in Figure 5(a), flooding the GaN/AlGaN interface with electrons that form a 2DEG channel. The output current flowing from the drain of the HEMT is then injected into the cathode of the LED, forward biasing it to the energy band diagram shown and indicated by the split quasi-Fermi levels of electrons $E_{fn}$ and holes $E_{fp}$. The electrons injected from the cathode radiatively recombine with the holes injected from the anode into the In0.17Ga0.83N quantum well, emitting light that is sensed by the photodiode.
Figure 5(c) shows that the photovoltage signal follows the gate voltage as it is swept between $-3.0\leq V_{\mathrm{GS}}\leq 3.0$ V with the drain voltage set at a voltage of 5.9 V and the backgate voltage set at a voltage of 1 V, over millisecond timescale across its threshold voltage from on state to off state. Figure 5(d) shows that when the gate voltage of the HEMT is modulated between $2.0\leq V_{\mathrm{GS}}\leq 3.0$ V when it is in its on state with $V_{\mathrm{GS}}>V_{\mathrm{T}}$, the light emitted by the LED follows this gate modulation, translating the electronic signal in the gate of the HEMT into a photonic bright and dim signal output from the LED on the other side of the wafer.
Due to the dualtronic structure, the cathode of the LED can also serve as a back-gate for the HEMT, which was taken into account in the monolithic switching measurements. With a separate contact to the n-GaN substrate we observed that the cathode voltage could exponentially control the drain current when the top gate was electrically floating. This back gating effect is available as a new functionality, or may be eliminated if undesirable by replacing the conductive substrate with a semi-insulating substrate. The backgating effect of the HEMT with the LED is shown in Extended Data Fig. 7.
The observations described here thus prove that the concept of dualtronics is feasible, opening the gates for many interesting possibilities. This work realized an electronics fabric on the N-polar face and a photonic fabric on the metal-polar face of the same wafer. This dualtronic combination is of immediate interest for combining microLEDs on the photonic side with transparent TFTs on the same GaN wafer on the other side. This monolithic convergence of devices reduces the number of components required for microLEDs with the potential for dramatic area and cost savings due to efficient use of the substrate real-estate. Moreover, monolithic integrating schemes combining HEMTs and LEDs on a single substrate face rely on multiple epitaxial layers that need to be either selectively removed and/or regrown in order to expose the LED heterostructure buried underneath the HEMT structure or vise versa [23, 24, 25, 26, 27]. Exposing the buried heterostructure by dry etching induces plasma damage, leading to diminished light emission and degradation of contacts. Similarly, with a selective growth method, the ex-situ nature of regrowth results in a poor growth interface, increasing leakage pathways. These issues are all avoided with the dualtronics integration scheme.
But the dualtronics concept extends to several exciting new opportunities. Laser diodes, semiconductor optical amplifiers (SOAs), and electro-optical modulators can be realized on one side of the wafer, and transistors on the other side, dramatically decreasing the components and chips desired in photonic integrated circuits. The combination of photodetectors and LEDs or Lasers on both sides can enable an on-chip photonic communication and computation system. RF transistor power amplifiers (PAs) for transmit part of communication systems can be realized on one polarity, and low-noise amplifiers (LNAs) for the receive end of communication systems can enable integrated transceivers in a combined transmit/receive module of smaller form factor than existing systems. The combination of n-channel transistors on one polarity with p-channel transistors on the opposite polarity can enable new forms of complementary transistor circuit topologies connected by through-vias through the substrate. Such dualtronic devices can take advantage of the wide bandgap nature of the polar nitride semiconductors for new forms of power electronics and RF electronics. These, and several allied possibilities can allow for creation and manipulation of electrons and photons on the opposite faces of the same wafer to achieve new functionalities.
The ultra wide bandgap polar semiconductor AlN boasts large electro-acoustic coupling, which makes it the preferred material today for acoustic wave RF filters. Dualtronics can thus take advantage of this property of the polar nitride semiconductors to combine sonar (by sound waves), radar (microwaves), and lidar (light) on the same platform. The efficient use of the substrate surfaces eliminates wasted space, reduces the energy and material costs of producing multiple wafers, and thus should be of high interest for future technologies well beyond the particular semiconductor materials discussed here.
Acknowledgements
Len van Deurzen and Eungkyun Kim contributed equally to this work.
The work at the Institute of High Pressure Physics PAS was supported by the National Science Centre (NCN) of Poland, Grant No. 2021/43/D/ST3/03266 and European Union under HORIZON EUROPE Digital, Industry and Space (VISSION ID:101070622),
This work was performed in part at the Cornell NanoScale Facility, a member of the National Nanotechnology Coordinated Infrastructure (NNCI), which is supported by the National Science Foundation (Grant NNCI-2025233). This work further made use of the electron microscopy facility of the Cornell Center for Materials Research (CCMR) with support from the National Science Foundation Materials Research Science and Engineering Centers (MRSEC) program (DMR1719875). The Thermo Fisher Spectra 300 X-CFEG was acquired with support from PARADIM, an NSF MIP (DMR-2039380) and Cornell University.